# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ Abstract
We study contextual bandits with low-rank structure where, in each round, if the (context, arm) pair $(i,j)\in[m]\times[n]$ is selected, the learner observes a noisy sample of the $(i,j)$-th entry of an unknown low-rank reward matrix. Successive contexts are generated randomly in an i.i.d. manner and are revealed to the learner. For such bandits, we present efficient algorithms for policy evaluation, best policy identification and regret minimization. For policy evaluation and best policy identification, we show that our algorithms are nearly minimax optimal. For instance, the number of samples required to return an $\varepsilon$-optimal policy with probability at least $1-\delta$ typically scales as ${m+n\over\varepsilon^{2}}\log(1/\delta)$. Our regret minimization algorithm enjoys minimax guarantees scaling as $r^{7/4}(m+n)^{3/4}\sqrt{T}$, which improves over existing algorithms. All the proposed algorithms consist of two phases: they first leverage spectral methods to estimate the left and right singular subspaces of the low-rank reward matrix. We show that these estimates enjoy tight error guarantees in the two-to-infinity norm. This in turn allows us to reformulate our problems as a misspecified linear bandit problem with dimension roughly $r(m+n)$ and misspecification controlled by the subspace recovery error, as well as to design the second phase of our algorithms efficiently.
@@ 1 Introduction
(b) Policy Evaluation (PE). For this task, the objective is to estimate the average reward of a given target policy based on data generated from a fixed behavior policy. We derive instance-specific and minimax lower bounds on the sample complexity satisfied by any $(\varepsilon,\delta)$-PAC algorithms [^1] These algorithms return an $\varepsilon$-accurate estimate of the value of the policy with certainty level $1-\delta$.. The latter typically [^2] By ”typically” here, we mean when the matrix $M$ and the (context,arm) distribution are homogeneous (refer to Definition 3.1 for details). scales as ${(m+n)\over\varepsilon^{2}}\log(1/\delta)$. We leverage our method to devise SIPS (Spectral Importance Propensity Score) and RS-PE (Recover Subspace for Policy Evaluation), two PE algorithms with nearly minimax optimal sample complexity. The second phase of RS-PE refines the estimate of the reward matrix, using a regularized least-square-estimator applied to the misspecified linear bandit model obtained in the first phase.
(c) Best Policy Identification (BPI). Here, the goal is to return an approximately optimal policy based on data gathered using a fixed or adaptive sampling strategy. Using the same estimates of the reward matrix as those used in the PE algorithms, we devise SBPI (Spectral Best Policy Identification) and RS-BPI (RS for Best Policy Identification), two $(\varepsilon,\delta)$-PAC algorithms with nearly minimax optimal sample complexity, again typically scaling as ${(m+n)\over\varepsilon^{2}}\log(1/\delta)$. This significantly improves over existing algorithms whose sample complexity scales as ${(m+n)\over\varepsilon^{2+r}}\log(1/\delta)$ (Lee et al., 2023).
(d) Regret minimization. The two-phase design method yields RS-RMIN (RS for Regret MINimization), an algorithm with minimax regret guarantees scaling as $(m+n)^{3/4}\sqrt{T}$ over $T$ rounds. Surprisingly and to the best of our knowledge, RS-RMIN is the first algorithm enjoying minimax guarantees that are always strictly tighter than those achieved in unstructured bandits (in this case, the best guarantees scale as $\sqrt{mnT}$). In its second phase, RS-RMIN uses an extension of SupLin-UCB (Chu et al., 2011) to solve the misspecified linear bandit problem derived at the end of the first phase.
@@
For a given matrix $M\in\mathbb{R}^{m\times n}$, we denote its $i$-th row by $M_{i,:}$, its $j$-th column by $M_{:,j}$, and its $(i,j)$-th entry by $M_{i,j}$. We denote $\|M\|_{\textup{op}}$ the Euclidean operator norm, $\|M\|_{\textup{F}}$ the Frobenius norm, $\|M\|_{2\to\infty}=\max_{i\in[m]}\|M_{i,:}\|_{2}$ the two-to-infinity norm, $\|M\|_{\max}=\max_{(i,j)\in[m]\times[n]}|M_{i,j}|$, $\lambda_{\max}\left(M\right)$ (resp. $\lambda_{\min}\left(M\right)$) the maximal (resp. minimal) eigenvalue. For a given vector $x$, we denote its Euclidean norm by $\|x\|_{2}$, and by $\|x\|_{\Lambda}=\sqrt{x^{\top}\Lambda x}$ its Euclidean norm weighted by some positive definite matrix $\Lambda$. The notation $f(x)\lesssim g(x)$ (resp. $f(x)\gtrsim g(x)$) means that there exists a universal constant $C>0$ such that $f(x)\leq Cg(x)$ (resp. $f(x)\geq Cg(x)$) for all $x$. We write $f(x)=\Theta\left(g(x)\right)$ when $f(x)\lesssim g(x)$ and $f(x)\gtrsim g(x)$. We write $\operatorname{poly}\left(x\right)$ to denote a quantity that is upper bounded by a polynomial function of $x$. We also use $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$. Finally, for any $p,q\in(0,1)$, we define $\operatorname{kl}(p,q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$ as the KL-divergence between two Bernoulli distributions with mean $p$ and $q$, respectively.
@@ 2 Related Work
There has been some interest in settings similar to ours (Gentile et al., 2014; Gopalan et al., 2016; Pal & Jain, 2022; Lee et al., 2023), although mostly in the context of regret minimization. Nonetheless, progress on minimax guarantees has remained surprisingly limited. For example, (Pal & Jain, 2022) proposed OCTAL, an algorithm that achieves a regret of order $O(\textrm{polylog}(m+n)\sqrt{T})$ for rank $1$ bandits, and a simple ETC-based algorithm achieving regret $O(\textrm{polylog}(m+n)T^{3/4})$ for general rank $r$. However, their results assume that the learner observes $m$ entries per round, while in our setting, only one entry is observed. Recently, (Lee et al., 2023) considered a setting closely related to ours, and referred to as context-lumpable bandit. There, the rows of the reward matrix can be clustered into $r$ groups, where within each group, the rows are identical. They establish minimax regret upper bounds of order $\widetilde{O}(\sqrt{r^{3}(m+n)T})$. They further extend their results to contextual low-rank bandits, but for these bandits, prove a regret upper bound of order $\widetilde{O}((m+n)^{\frac{1}{3r+2}}T^{\frac{3r+1}{3r+2}})$ (see Theorem 26 in (Lee et al., 2023)). They leave the existence of an algorithm with minimax regret of order $\widetilde{O}(\sqrt{\textrm{poly}(r)(m+n)T})$ as an open question. We do not fully answer this question, but propose an algorithm with regret scaling as $\widetilde{O}(r^{7/4}(m+n)^{3/4}\sqrt{T})$, which improves over existing algorithms.
@@
Singular subspace recovery in the two-to-infinity norm. It is not surprising that estimating the singular subspaces of $M$ is useful in the context of bandit problems with low-rank structure as it has been showcased in (Jun et al., 2019; Lale et al., 2019; Lu et al., 2021; Kang et al., 2022). However, to the best of our knowledge, only Frobenius norm guarantees have been used, and guarantees in the two-to-infinity norm have remained largely unexplored. The latter guarantees are often harder to obtain and progress towards obtaining them have only emerged recently (Eldridge et al., 2018; Fan et al., 2018; Cape et al., 2019; Abbe et al., 2020). Typically, under standard assumptions, such as bounded incoherence (Candès & Tao, 2010; Recht, 2011), these guarantees state that, for a matrix of size $m\times n$, the subspace estimation error in the two-to-infinity norm is at least smaller by a factor of $\sqrt{1/(m+n)}$ than the achievable bounds on the estimation error in the Frobenius norm. For this reason, we say that the subspace recovery error exhibits a delocalization phenomenon (Rudelson & Vershynin, 2015), i.e., the estimation error is spread out across $m+n$ directions (see also the survey (Chen et al., 2021) and references therein).
The major difficulty towards using two-to-infinity norm guarantees is that their derivation requires stringent independence assumptions. Recently, (Stojanovic et al., 2023) provided tools to relax such assumptions and to accommodate scenarios that are suitable for sequential decision making problems such as reinforcement learning and bandits. We make use of these tools to obtain tight guarantees on the singular subspace recovery in the two-to-infinity norm for low-rank bandits. These guarantees enable an effective reduction of the contextual low-rank bandit problem to the misspecified contextual linear bandit problem.
@@ 3 Model and Learning Objectives
We consider stochastic bandit problems with low-rank structure. Specifically, we assume that the expected rewards are parametrized by a matrix $M\in\mathbb{R}^{m\times n}$ which has low rank $r\ll\min(m,n)$. Its SVD is $M=U\Sigma V^{\top}$, where $U\in\mathbb{R}^{m\times r}$ (resp. $V\in\mathbb{R}^{n\times r}$) contains the left (resp. right) singular vectors, and where $\Sigma=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})$ and $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}$. The incoherence parameters of $M$ are defined as $\mu(U)=\sqrt{m/r}\|U\|_{2\to\infty}$ and $\mu(V)=\sqrt{n/r}\|V\|_{2\to\infty}$. We define $\mu=\max\left\{\mu(U),\mu(V)\right\}$, and denote by $\kappa=\sigma_{1}/\sigma_{r}$ the condition number of $M$.
Contextual low-rank bandits. In each round $t\geq 1$, the learner first observes a context $i_{t}$, selected in an i.i.d. manner and with distribution $\rho$ over $[m]$. She then selects an arm $j_{t}\in[n]$ potentially based on previous observations, and receives a reward $r_{t}=M_{i_{t},j_{t}}+\xi_{t}$. The reward matrix $M$ is a priori unknown with $\|M\|_{\max}=\Theta(1)$, and $(\xi_{t})_{t\geq 1}$ is a sequence of i.i.d. zero-mean and $1$-subgaussian random variables. We denote the context-arm distribution by $\omega.$
For a given randomized policy $\pi$, we define its policy value as $v^{\pi}:=\sum_{i,j}\omega^{\pi}_{i,j}M_{i,j}$ with $\omega^{\pi}_{i,j}:=\rho_{i}\pi\left(j|i\right)$ ($\pi(j|i)$ is the probability to play arm $j$ when observing context $i$). We further define the optimal policy $\pi^{\star}:=\operatorname*{arg\,max}_{\pi}v^{\pi}$ and denote its value by $v^{\star}$. We consider three learning tasks:
@@
    Policy evaluation. Given a target policy $\pi$, and assuming data is gathered under a policy [^3] The arm $j$ is selected with probability $\pi^{b}(j|i)$ when the context $i$ is observed. $\pi^{b}$, we aim at designing an efficient estimator of its policy value $v^{\pi}$. We say that a PE estimator $\hat{v}^{\pi}$ is $(\varepsilon,\delta)$-PAC if $\mathbb{P}_{M}\left(|v^{\pi}-\widehat{v}^{\pi}|\leq\varepsilon\right)\geq 1-\delta$ for every rank-$r$ matrix $M$.
@@
    Best policy identification. We aim to design an efficient algorithm to identify an $\varepsilon$-optimal policy. We say that a BPI algorithm is $\left(\varepsilon,\delta\right)$-PAC if it outputs a policy $\hat{\pi}:[m]\to[n]$ such that $\mathbb{P}_{M}\left(v^{\star}-v^{\hat{\pi}}\leq\varepsilon\right)\geq 1-\delta,$ for every $M\in\mathbb{R}^{m\times n}$ of rank $r$.
@@
The reward matrix $M$ is homogeneous when $m=\Theta(n)$, $r=\Theta(1)$, $\kappa=\Theta(1)$, $\mu=\Theta(1)$. A distribution $p$ on a finite set ${\cal I}$ is homogeneous when $p_{\min}=\Theta\left(p_{\max}\right)$ for $p_{\min}=\min_{i\in{\cal I}}p_{i}$, $p_{\max}=\max_{i\in{\cal I}}p_{i}$.
The notion of homogeneity allows us to have meaningful discussions but also constitutes a reasonable assumption. In particular, assuming that the matrix $M$ has bounded incoherence entails that the matrix $M$ can be recovered using fewer samples than $mn$ (Candes & Recht, 2012). The bounded incoherence property is also related to the notion of spikiness of a matrix, which has to do with assuming that $\|M\|_{\max}/\sigma_{1}(M)=\Theta(1/\sqrt{mn})$ (Negahban & Wainwright, 2012; Mackey et al., 2015)). Indeed, we can verify (see Lemma B.8) that $\frac{1}{\sqrt{mn}}\leq\frac{\|M\|_{\max}}{\sigma_{1}(M)}\leq\frac{\mu^{2}r}{\sqrt{mn}}$. Note that this inequality does not mean that $\|M\|_{\max}=\Theta(1/\sqrt{mn})$ when $\mu=\Theta(1)$, because for most interesting settings $\sigma_{1}(M)$ would scale as $\sqrt{mn}$. A naive example would be to take a matrix $M$ that has all entries equal to $1$. Observe then that $\textup{rank}(M)=1$, $\|M\|_{\max}=1$, $\sigma_{1}(M)=\sqrt{mn}$, and $M$ satisfies the bounded incoherence assumption.

### v2

@@ Abstract
We study contextual bandits with low-rank structure where, in each round, if the (context, arm) pair $(i,j)\in[m]\times[n]$ is selected, the learner observes a noisy sample of the $(i,j)$-th entry of an unknown low-rank reward matrix. Successive contexts are generated randomly in an i.i.d. manner and are revealed to the learner. For such bandits, we present efficient algorithms for policy evaluation, best policy identification and regret minimization. For policy evaluation and best policy identification, we show that our algorithms are nearly minimax optimal. For instance, the number of samples required to return an $\varepsilon$-optimal policy with probability at least $1-\delta$ typically1 scales as ${r(m+n)\over\varepsilon^{2}}\log(1/\delta)$. Our regret minimization algorithm enjoys minimax guarantees typically [^1] By ”typically” here, we mean when the matrix $M$ and the (context, arm) distribution are homogeneous (refer to Definition 3.1 for details). Note that under homogeneity of $M$, $m=\Theta(n)$. scaling as $r^{5/4}(m+n)^{3/4}\sqrt{T}$, which improves over existing algorithms. All the proposed algorithms consist of two phases: they first leverage spectral methods to estimate the left and right singular subspaces of the low-rank reward matrix. We show that these estimates enjoy tight error guarantees in the two-to-infinity norm. This in turn allows us to reformulate our problems as a misspecified linear bandit problem with dimension roughly $r(m+n)$ and misspecification controlled by the subspace recovery error, as well as to design the second phase of our algorithms efficiently.
@@ 1 Introduction
(b) Policy Evaluation (PE). For this task, the objective is to estimate the average reward of a given target policy based on data generated from a fixed behavior policy. We derive instance-specific and minimax lower bounds on the sample complexity satisfied by any $(\varepsilon,\delta)$-PAC algorithms [^2] These algorithms return an $\varepsilon$-accurate estimate of the value of the policy with certainty level $1-\delta$.. The latter typically1 scales as ${m+n\over\varepsilon^{2}}\log(1/\delta)$. We leverage our method to devise SIPS (Spectral Importance Propensity Score) and RS-PE (Recover Subspace for Policy Evaluation), two PE algorithms with nearly minimax optimal sample complexity. The second phase of RS-PE refines the estimate of the reward matrix, using a regularized least-square-estimator applied to the misspecified linear bandit model obtained in the first phase.
(c) Best Policy Identification (BPI). Here, the goal is to return an approximately optimal policy based on data gathered using a fixed or adaptive sampling strategy. Using the same estimates of the reward matrix as those used in the PE algorithms, we devise SBPI (Spectral Best Policy Identification) and RS-BPI (RS for Best Policy Identification), two $(\varepsilon,\delta)$-PAC algorithms with nearly minimax optimal sample complexity, again typically scaling as ${m+n\over\varepsilon^{2}}\log(1/\delta)$. This significantly improves over existing algorithms whose sample complexity scales as ${m+n\over\varepsilon^{2+r}}\log(1/\delta)$ (Lee et al., 2023).
(d) Regret minimization. The two-phase design method yields RS-RMIN (RS for Regret MINimization), an algorithm with minimax regret guarantees typically scaling as $(m+n)^{3/4}\sqrt{T}$ over $T$ rounds. Surprisingly and to the best of our knowledge, RS-RMIN is the first algorithm enjoying minimax guarantees that are always strictly tighter than those achieved in unstructured bandits (in this case, the best guarantees scale as $\sqrt{mnT}$). In its second phase, RS-RMIN uses an extension of SupLinUCB (Chu et al., 2011; Takemura et al., 2021) to solve the misspecified linear bandit problem derived at the end of the first phase.
@@
For a given matrix $M\in\mathbb{R}^{m\times n}$, we denote its $i$-th row by $M_{i,:}$, its $j$-th column by $M_{:,j}$, and its $(i,j)$-th entry by $M_{i,j}$. We denote by $\|M\|_{\textup{op}}$ its Euclidean operator norm, by $\|M\|_{\textup{F}}$ its Frobenius norm, by $\|M\|_{2\to\infty}=\max_{i\in[m]}\|M_{i,:}\|_{2}$ its two-to-infinity norm, by $\|M\|_{\max}=\max_{(i,j)\in[m]\times[n]}|M_{i,j}|$ its max-norm, by $\lambda_{\max}\left(M\right)$ (resp. $\lambda_{\min}\left(M\right)$) its maximal (resp. minimal) eigenvalue. For a given vector $x$, we denote its Euclidean norm by $\|x\|_{2}$, and by $\|x\|_{\Lambda}=\sqrt{x^{\top}\Lambda x}$ its Euclidean norm weighted by some positive definite matrix $\Lambda$. The notation $f(x)\lesssim g(x)$ (resp. $f(x)\gtrsim g(x)$) means that there exists a universal constant $C>0$ such that $f(x)\leq Cg(x)$ (resp. $f(x)\geq Cg(x)$) for all $x$. We write $f(x)=\Theta\left(g(x)\right)$ when $f(x)\lesssim g(x)$ and $f(x)\gtrsim g(x)$. We write $\operatorname{poly}\left(x\right)$ to denote a quantity that is upper bounded by a polynomial function of $x$. We also use $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$. Finally, for any $p,q\in(0,1)$, we define $\operatorname{kl}(p,q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$ as the KL-divergence between two Bernoulli distributions with mean $p$ and $q$, respectively.
@@ 2 Related Work
There has been some interest in settings similar to ours (Gentile et al., 2014; Gopalan et al., 2016; Pal & Jain, 2022; Lee et al., 2023; Pal et al., 2023), although mostly in the context of regret minimization. Nonetheless, progress on minimax guarantees has remained surprisingly limited. For example, (Pal & Jain, 2022) proposed OCTAL, an algorithm that achieves a regret of order $O(\mathrm{polylog}(m+n)\sqrt{T})$ for rank-$1$ bandits, and a simple ETC-based algorithm achieving a regret of $O(\mathrm{polylog}(m+n)T^{3/4})$ for general rank $r$. However, their results assume that the learner observes $m$ entries per round, while in our setting, only one entry is observed. Recently, (Lee et al., 2023) considered a setting closely related to ours, and referred to as context-lumpable bandits. There, the rows of the reward matrix can be clustered into $r$ groups, within which the rows are identical. They establish minimax regret upper bounds of order $\widetilde{O}(\sqrt{r^{3}(m+n)T})$. They further extend their results to contextual low-rank bandits, but for these bandits, prove a regret upper bound of order $\widetilde{O}((m+n)^{\frac{1}{3r+2}}T^{\frac{3r+1}{3r+2}})$ (see Theorem 26 in (Lee et al., 2023)). A cluster-like structure comparable to that of (Lee et al., 2023) was also considered in (Pal et al., 2023), where the authors proposed an algorithm that attains a regret of order $\tilde{O}(\sqrt{\mathrm{poly}(r)(m+n)T})$. These works leave the existence of an algorithm with minimax regret of order $\widetilde{O}(\sqrt{\mathrm{poly}(r)(m+n)T})$ in contextual low-rank bandit as an open question. We do not fully answer this question, but propose an algorithm with regret scaling as $\widetilde{O}(r^{5/4}(m+n)^{3/4}\sqrt{T})$, which improves over existing algorithms.
@@
Singular subspace recovery in the two-to-infinity norm. It is not surprising that estimating the singular subspaces of $M$ is useful in the context of bandit problems with low-rank structure as it has been showcased in (Jun et al., 2019; Lale et al., 2019; Lu et al., 2021; Kang et al., 2022). However, to the best of our knowledge, only Frobenius norm guarantees have been used, and guarantees in the two-to-infinity norm have remained largely unexplored. The latter guarantees are often harder to obtain and progress towards obtaining them have only emerged recently (Eldridge et al., 2018; Fan et al., 2018; Cape et al., 2019; Abbe et al., 2020). Typically, under standard assumptions, such as bounded incoherence (Candès & Tao, 2010; Recht, 2011), these guarantees state that, for a matrix of size $m\times n$, the subspace estimation error in the two-to-infinity norm is smaller by a factor of $\sqrt{m+n}$ than the achievable bounds on the estimation error in the Frobenius norm. For this reason, we say that the subspace recovery error exhibits a delocalization phenomenon (Rudelson & Vershynin, 2015), i.e., the estimation error is spread out across $m+n$ directions (see also the survey (Chen et al., 2021) and references therein).
The major difficulty towards using two-to-infinity norm guarantees is that their derivation requires stringent independence assumptions. Recently, (Stojanovic et al., 2024) provided tools to relax such assumptions and to accommodate scenarios that are suitable for sequential decision making problems such as Markov decision processes and bandits. We make use of these tools to obtain tight guarantees on the singular subspace recovery in the two-to-infinity norm for low-rank bandits. These guarantees enable an effective reduction of the contextual low-rank bandit problem to the misspecified contextual linear bandit problem.
@@ 3 Model and Learning Objectives
We consider a stochastic bandit problem with low-rank structure. Specifically, we assume that the expected rewards are parametrized by a matrix $M\in\mathbb{R}^{m\times n}$ which has low rank $r\ll\min(m,n)$. Its SVD is $M=U\Sigma V^{\top}$, where $U\in\mathbb{R}^{m\times r}$ (resp. $V\in\mathbb{R}^{n\times r}$) contains the left (resp. right) singular vectors, and where $\Sigma=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})$ with $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{r}$. The incoherence parameters of $M$ are defined as $\mu(U)=\sqrt{m/r}\|U\|_{2\to\infty}$ and $\mu(V)=\sqrt{n/r}\|V\|_{2\to\infty}$. We define $\mu=\mu(U)\vee\mu(V)$, and denote by $\kappa=\sigma_{1}/\sigma_{r}$ the condition number of $M$.
Contextual low-rank bandits. In each round $t\geq 1$, the learner first observes a context $i_{t}$, selected in an i.i.d. manner and with distribution $\rho$ over $[m]$. She then selects an arm $j_{t}\in[n]$ potentially based on previous observations, and receives a reward $r_{t}=M_{i_{t},j_{t}}+\xi_{t}$. The reward matrix $M$ is a priori unknown, and $(\xi_{t})_{t\geq 1}$ is a sequence of i.i.d. zero-mean and $\sigma$-subgaussian random variables. We denote the (context, arm) distribution by $\omega.$
For a given randomized policy $\pi$, we define its policy value as $v^{\pi}:=\sum_{i,j}\omega^{\pi}_{i,j}M_{i,j}$ with $\omega^{\pi}_{i,j}:=\rho_{i}\pi\left(j|i\right)$ ($\pi(j|i)$ is the probability to select arm $j$ when observing context $i$). We further define the optimal policy $\pi^{\star}:=\operatorname*{arg\,max}_{\pi}v^{\pi}$ and denote its value by $v^{\star}$. We consider three learning tasks:
@@
    Policy evaluation. Given a target policy $\pi$, and assuming data is gathered under a policy [^3] The arm $j$ is selected with probability $\pi^{b}(j|i)$ when the context $i$ is observed. $\pi^{b}$, we aim at designing an efficient estimator of its policy value $v^{\pi}$. We say that a PE estimator $\hat{v}^{\pi}$ is $(\varepsilon,\delta)$-PAC if $\mathbb{P}_{M}\left(|v^{\pi}-\hat{v}^{\pi}|\leq\varepsilon\right)\geq 1-\delta$ for every rank-$r$ matrix $M\in\mathbb{R}^{m\times n}$.
@@
    Best policy identification. We aim to design an efficient algorithm to identify an $\varepsilon$-optimal policy. We say that a BPI algorithm is $\left(\varepsilon,\delta\right)$-PAC if it outputs a policy $\hat{\pi}:[m]\to[n]$ such that $\mathbb{P}_{M}\left(v^{\star}-v^{\hat{\pi}}\leq\varepsilon\right)\geq 1-\delta$ for every rank-$r$ matrix $M\in\mathbb{R}^{m\times n}$.
@@
The reward matrix $M$ is homogeneous when $m=\Theta(n)$, $r=\Theta(1)$, $\mu=\Theta(1)$, $\kappa=\Theta(1)$. A distribution $p$ on a finite set ${\cal I}$ is homogeneous when $p_{\min}=\Theta\left(p_{\max}\right)$ for $p_{\min}=\min_{i\in{\cal I}}p_{i}$, $p_{\max}=\max_{i\in{\cal I}}p_{i}$  [^4] This implies $p_{\min}=\Theta\left(1/|\cal I|\right)$, $p_{\max}=\Theta\left(1/|\cal I|\right)$. .
The notion of homogeneity allows us to have meaningful discussions but also constitutes a reasonable assumption. In particular, assuming that the matrix $M$ has bounded incoherence entails that the matrix $M$ can be recovered using fewer samples than $mn$ (Candes & Recht, 2012). The bounded incoherence property is also related to the notion of spikiness of a matrix, which has to do with assuming that $\|M\|_{\max}/\sigma_{1}=\Theta(1/\sqrt{mn})$ (Negahban & Wainwright, 2012; Mackey et al., 2015)). Indeed, we can verify (see Lemma B.8) that $\frac{1}{\sqrt{mn}}\leq\frac{\|M\|_{\max}}{\sigma_{1}}\leq\frac{\mu^{2}r}{\sqrt{mn}}$. Note that this inequality does not mean that $\|M\|_{\max}=\Theta(1/\sqrt{mn})$ when $\mu=\Theta(1)$, because for most interesting settings $\sigma_{1}$ would scale as $\sqrt{mn}$. A naive example would be to take a matrix $M$ that has all entries equal to $1$. Observe then that $\textup{rank}(M)=1$, $\|M\|_{\max}=1$, $\sigma_{1}=\sqrt{mn}$, and $M$ satisfies the bounded incoherence assumption.